# Adaptation of Agentic AI

### v2

@@ 4.1 A1: Tool Execution Result as Signal
Tool execution result as signal refers to a class of adaptation mechanisms where an agent leverages the actual outcomes of external tool invocations (such as execution correctness, functional success, or numerical improvement) as feedback to refine its behavior. Here, the tool or environment serves as an objective source of feedback, forming a verifiable signal that can drive both supervised and reinforcement-based learning.
@@ 4.1.1 Earlier Works: SFT & Off-Policy Methods
Early A1-type methods typically focus on SFT or DPO, which aim to teach agents using pre-collected data. These methods begin by collecting a set of model responses or trajectories involving tool usage, and then use this data for supervised fine-tuning or DPO. These A1 methods share a common foundation in leveraging objective environment-grounded outcomes, but differ in the form, source, and utilization of their feedback signals. The evolution of these imitation-based A1-type methods primarily centers on the transformation in how training signal is obtained and utilized.
The earliest representative, Toolformer [4] (NeurIPS 2023), introduced the idea of using tool outcomes as self-supervised learning signals. The model automatically inserts candidate API calls into text, executes them, and measures whether the returned result improves token prediction likelihood. A call is retained if it significantly reduces perplexity, formalized as $L_{i}^{-}-L_{i}^{+}\geq\tau_{f}$, where the reduction quantifies a self-supervised tool execution result signal. This implicit signal anchors learning in the correctness of tool usage, enabling the model to autonomously discover when external APIs improve performance.
However, since the training is based on self-supervised feedback, Toolformer remains limited in precision when applied to real executable environments. This limitation motivated a series of subsequent approaches that sought to introduce more reliable, externally grounded learning signals. Building upon this insight, the evolution of A1-type methods can be viewed through three progressively grounded paradigms of learning:
  • Alignment with golden answers: supervision comes from correct responses or expert trajectories.
  • Alignment with golden formats: correctness is defined structurally or syntactically.
  • Alignment with direct tool execution: learn from verifiable outcomes produced by executing tools, allowing supervision to emerge from actual tool behavior rather than predefined labels.
Alignment with golden answers. Early A1-type approaches focused on aligning models with correct final outputs, typically defined by task-specific ground truths or verified expert solutions.
TRICE [52] (NAACL 2024) is a two-stage framework designed to teach LLMs when and how to use tools selectively. The first stage uses supervised fine-tuning to provide the model with a preliminary ability to imitate tool-use behavior. The core of the method lies in the second stage, “Reinforcement Learning with Execution Feedback (RLEF)”. In this stage, the agent is trained using a reward signal derived directly from tool execution. The system collects a set of candidate responses for a given task, some of which involve tool calls. A reward strategy then scores each response by comparing its execution result against the ground-truth answer. The model is then reinforced, using a ranking loss, to align its preferences with the high-reward responses, effectively learning from the successful or failed outcomes of tool execution to mitigate excessive reliance on tools and improve accuracy. This design provides a clear instance of learning grounded in correctness alignment, where rewards are explicitly tied to whether the tool-executed outcome matches the golden answer.
ToolAlpaca [53] represents one of the earliest closed-loop implementations of A1-type adaptation, where the model refines its tool-use capability directly through iterative interaction with executable environments. The model first generates a tool-call candidate, for example an API invocation, which is then executed within the environment. The system records the runtime outcome of each execution, including returned values, errors, or completion states, to determine the correctness of the model’s action. This design establishes an automated self-improvement loop consisting of four key stages: generate, execute, evaluate, and finetune. Through repeated cycles, ToolAlpaca progressively aligns its internal representation with the actual semantics and behavior of the tools it uses. This closed-loop process embodies the essence of the A1 paradigm, where tool execution results themselves serve as the primary adaptation signal. ToolAlpaca demonstrates the ability to generalize across unseen tools and to adapt its calling strategy according to contextual requirements. By grounding updates in correctness relative to observed outcomes, it implicitly aligns its learning with the “golden answer” paradigm.
TP-LLaMA [54] (NeurIPS 2024) is an inference trajectory optimization framework designed to improve tool-augmented LLMs by learning from errors. The authors observe that prior models like ToolLLaMA [10] are trained via SFT exclusively on successful expert trajectories from the ToolBench dataset, which ignores valuable information contained in failed exploration paths of the decision trees. To address this, their method consists of two stages: first, the model undergoes standard SFT on successful trajectories, similar to the baseline. Second, the framework leverages the "failed paths" by constructing a novel preference dataset called ToolPreference. This dataset is built using a step-wise method: for any decision node along a successful path, the "preferred" output ($y_{w}$) is the expert’s correct next step, while the dispreferred output ($y_{l}$) is any corresponding failed branch originating from that same node. The model is then trained on these preference pairs using DPO. This approach explicitly uses the execution feedback from failed attempts as a training signal, enabling the model to learn from failure and significantly enhancing its decision-making, generalization, and reasoning efficiency. TP-LLaMA therefore transforms failure signals into preference-aligned supervision, reinforcing learning according to correctness with respect to expert trajectories.
Alignment with golden formats. A complementary branch of A1-type evolution shifted from output-based correctness to format-based structural correctness, emphasizing syntactic and logical validity of tool calls rather than explicit task answers.
Gorilla [55] (NeurIPS 2024) is a retrieval-augmented LLaMA-based language model fine-tuned to generate correct API calls across a large and changing set of machine learning APIs. During training, it leverages self-instructed instruction–API pairs and, optionally, a document retriever to adapt to test-time documentation changes. A crucial component of Gorilla’s evaluation and feedback loop is the use of Abstract Syntax Trees (ASTs). Both the model-generated API calls and reference API calls are converted into ASTs, and correctness is determined by checking whether the reference API forms a subtree of the generated AST. Compared to direct text matching, AST-based evaluation is more robust: differences in parameter order or optional arguments do not lead to false negatives, as ASTs focus on the logical structure of the API call. In the context of A1-type adaptation, this serves as a form of tool execution result signal, providing the model with structured feedback on whether its API call was functionally correct, which can then be used to guide learning and improve tool-use performance. Gorilla therefore exemplifies the golden-format paradigm, where correctness is defined by adherence to canonical structural representations rather than output values.
ToolFlow [56] (NAACL 2025) is a data synthesis pipeline designed to enhance the tool-calling capabilities of LLMs through the generation of natural and coherent dialogues. Traditional SFT approaches rely on synthetically generated tool-call data. However, previous methods often suffer from low diversity and limited coherence because tools are sampled randomly and dialogues are synthesized as single-turn interactions, ignoring multi-turn dependencies. ToolFlow addresses these limitations by introducing two key strategies: Graph-based Sampling and Planned Generation. The Graph-based Sampling strategy constructs a tool graph based on parameter and return-value similarities between tools. Nodes represent tools, and edges indicate their relevance, allowing the selection of tool subsets that are likely to interact effectively. This facilitates the generation of complex user requirements involving multiple interconnected tools. The Planned Generation strategy enables the LLM to first create a high-level dialogue plan that organizes user requests across multiple turns, including both tool-call tasks and non-tool interactions. This ensures logical consistency and natural flow throughout the dialogue, resulting in more realistic training data. Overall, ToolFlow provides a systematic approach for generating high-quality multi-turn tool-call dialogues that closely reflect real-world interaction scenarios, and exemplifies structural alignment through graph-based format consistency.
Alignment with direct tool execution. The most advanced stage of A1 evolution centers on learning directly from verifiable environment signals, where tool execution outcomes themselves become the supervision source.
CodeAct [57] (ICML 2024) represented a paradigm in which LLMs learn tool use through direct interaction with executable code environments. Instead of producing textual or JSON-based commands, the model generates executable code actions that are run within a sandboxed environment. The environment returns explicit execution feedback—such as success or failure signals and resulting outputs—which are then used as supervision to refine the model. This process grounds the learning objective in the verifiable outcomes of tool execution, rather than in human-annotated correctness or model-predicted preferences. Through this execution-based feedback loop, CodeAct effectively aligns model behavior with the underlying causal mechanisms of tools.
NExT [58] (ICML 2024) is a method designed to teach LLMs to reason about code execution, specifically for program repair tasks. The approach utilizes an iterative self-training loop based on a “Sample–Filter–Train” process. The core of this method lies in its filtering step, which uses a tool’s execution result as the primary training signal. An external tool, takes the agent’s generated code output and validates it by running a set of unit tests. The binary pass-or-fail diagnostic from this execution serves as a reward signal. Only the candidate solutions (both rationale and code) that successfully pass all unit tests are deemed correct and are collected into a new synthetic dataset, which is then used to finetune the agent for the next iteration. This iterative refinement, guided strictly by the tool’s verification, progressively improves the model’s ability to generate accurate fixes and high-quality, execution-aware rationales.
ToolLLM [10] and AutoTools [59] (WWW 2025) push this paradigm toward autonomous tool learning. Specifically, AutoTools consists of two core stages: tool encapsulation and tool programming. In the tool encapsulation stage, the LLM automatically parses raw API documentation and transforms each tool into a callable function, including structured docstrings, argument specifications, and usage examples. Syntax and runtime correctness are verified through an integration verification procedure, which checks not only individual functions but also input–output dependencies among related tools. Verified functions are aggregated into a function library for subsequent use. In the tool programming stage, the LLM directly generates executable programs that sequentially call these functions, resolve intermediate outputs, and ultimately solve user queries. This design allows the model to flexibly integrate multiple tools using a unified programming language, instead of relying on specialized tokens or handcrafted formats. To further enhance model expertise, AutoTools introduces AutoTools-Learning, a multi-task learning approach that trains the LLM on three synthetic tasks: (1) documentation understanding, (2) relevance learning for selecting appropriate tools, and (3) function learning for generating correct multi-step tool programs. The training dataset comprises 34k high-quality examples synthesized from public sources and existing benchmarks, providing fine-grained supervision for both tool understanding and programmatic reasoning. Through this combination of automated tool encapsulation, programmatic integration, and multi-task training, AutoTools represents a significant step towards fully autonomous A1-type adaptation. Execution outcomes of each tool call serve as direct feedback for iterative self-improvement, enabling LLMs to handle complex, multi-step tool-use tasks without explicit human intervention.
LeReT [60] (ICLR 2025) is a reinforcement learning framework for improving multi-hop retrieval in LLM pipelines. A query-generating LLM $\pi_{r}$ produces diverse search queries using few-shot prompt ensembles, and retrieved documents are scored with a reward function $R$. Queries are converted into preference pairs $(y_{i},y_{j})$, where the higher-reward query is preferred. The model is then fine-tuned using preference-based reinforcement learning, specifically Identity Policy Optimization (IPO). IPO is a method for directly optimizing a model to reflect human or reward-based preferences. Instead of modeling the probability of a preferred output as in DPO, IPO explicitly enforces that the model’s implicit reward difference between preferred and dispreferred outputs matches a target margin. Formally, it minimizes the squared deviation of the reward difference from a fixed margin:
@@
where $\tilde{r}_{\phi}(x,y)=\log\frac{\pi_{\phi}(y|x)}{\pi_{\text{ref}}(y|x)}$, $y_{w}$ and $y_{l}$ denote the preferred and dispreferred queries, $x$ is the context, and $\tau$ is a margin hyperparameter controlling the target difference between rewards. Intuitively, IPO encourages the model to assign higher scores to better outputs and ensures the difference reaches a predefined magnitude, providing a direct and stable training signal. LeReT can be applied iteratively, using the fine-tuned model to generate better exploration data in subsequent iterations. This method improves both retrieval accuracy and downstream generation quality, and can adapt to arbitrary off-the-shelf retrievers without modifying the generator model. As such, it represents the culmination of environment-grounded A1 learning, where reward optimization is explicitly tied to real-world, verifiable signals. Similarly, RetPO [61] (NAACL 2025) trains a query reformulation model to translate LLM-generated queries into retriever-optimal queries. The training signal is ingenious in its simplicity: GPT-4 generates multiple candidate rewrites, each is evaluated by running it through an off-the-shelf retriever (e.g., BM25), and the retrieval performance (measured by how well the retrieved documents support the correct answer) serves as the reward. A smaller, open-source LM is then trained via DPO to produce high-reward rewrites.
Overall, the evolution of A1-type methods reflects a gradual shift from implicit, self-supervised feedback to explicit, execution-grounded learning signals. Early methods such as Toolformer demonstrated the feasibility of using tool execution as a self-supervised signal, but their reliance on internal likelihood metrics limited their external validity. Subsequent approaches strengthened the supervision source by grounding model optimization in correctness relative to golden answers or expert trajectories, allowing models to learn more directly from verified outcomes. The next wave of methods, such as Gorilla and ToolFlow, advanced this idea by emphasizing structural and syntactic alignment, ensuring that generated tool calls conform to canonical formats even under distributional shifts. Finally, environment-grounded methods fully integrated external feedback loops, enabling models to learn directly from verifiable execution signals and real-world rewards. This progression illustrates a clear trend toward deeper coupling between model reasoning and environment interaction, where feedback transitions from probabilistic imitation to causally grounded supervision. Such grounding not only enhances tool reliability and adaptability but also marks a key step toward autonomous, self-improving LLM agents capable of operating robustly in open, dynamic environments.
Despite these advances, these methods rely on SFT or DPO, where model updates are derived from pre-collected trajectories or candidate responses. While effective for exploiting existing data, these approaches are inherently constrained in exploration and may fail to fully capture the dynamic nature of interactive environments.
@@ 4.1.2 RLVR-Based Methods
Reinforcement learning with verifiable reward (RLVR) marks a pivotal stage in the evolution of A1-type adaptation, where models learn directly from online interaction with tools and environments. Unlike SFT or DPO approaches that rely on pre-collected trajectories or candidate responses, RLVR-based methods enable LLMs to iteratively explore, execute, and refine their actions based on immediate environment feedback. This paradigm allows adaptation to be dynamic, context-aware, and tightly coupled with the specific execution environment, spanning diverse domains such as web search, code generation, multi-tool reasoning, and downstream applications.
Web search and information retrieval tools optimized query generation and retrieval using environment-derived rewards. DeepRetrieval [21] (COLM 2025) marked a key turning point in A1-type adaptation, introducing RLVR to train LLMs as search agents that learn directly from retrieval outcomes. It formalizes query reformulation as an MDP where the user query is the state, the rewritten query is the action, and retrieval metrics—such as Recall@K, NDCG, or SQL execution accuracy—serve as the reward. The policy is optimized via KL-regularized PPO:
@@
where $r(q,q^{\prime})=r_{\text{retrieval}}(q,q^{\prime})+r_{\text{format}}(q^{\prime})$ jointly captures retrieval effectiveness and syntactic validity. This unified formulation enables the same framework to adapt seamlessly across literature search, QA-style retrieval, local corpus search with dense or BM25 retrievers, and text-to-SQL database querying. Empirically, DeepRetrieval achieved roughly a threefold improvement in recall (65.1% vs. previous state-of-the-art 24.7%) on literature search tasks using real-world search engines, while maintaining impressive performance across other retrieval and SQL domains. These results established reinforcement learning on environment rewards as a general, scalable, and cost-efficient paradigm for retrieval-based agent adaptation. ReZero [62], a successor to DeepRetrieval, extends this idea with GRPO-based reinforcement learning that rewards adaptive retries after failed searches. By introducing retry-aware reward shaping, it improves agent persistence and robustness in dynamic or partially observable web environments, further validating the effectiveness of DeepRetrieval’s reinforcement-driven approach. Orion [63] further extends DeepRetrieval by moving from single-step reformulation to multi-turn adaptive search. Using GRPO with turn-level rewards based on normalized similarity and rank, Orion trains models to iteratively refine, pivot, or backtrack through structured think-search cycles. This yields strong multi-hop retrieval performance with compact 350M–1.2B models, showing that effective multi-step search strategies can be learned without large controllers.
Code-based tools provided deterministic or sandboxed execution environments for reasoning and task completion. LeDex [64] (NeurIPS 2024) applied reinforcement learning using a PPO-based algorithm with a novel reward function that considers both the correctness of the refined code (via unit test results and CodeBLEU) and the quality of the explanation (via semantic similarity). RLEF [20] (ICML 2025) framed code synthesis as a multi-turn interactive task, where the LLM generates a solution, receives automatic feedback from executing the code on public test cases, and updates its subsequent generations accordingly. The process is formalized as a partially observable Markov Decision Process (MDP), with actions corresponding to token-level code generation and rewards determined by the success of private test cases, optimized using PPO. Code-R1 [65] built a reliable, scalable, and sandboxed reward pipeline to minimize reward false positives caused by faulty tests, unsolvable prompts, or mismatched execution environments. The key finding is that reward quality matters more than data quantity — clean, verified datasets and secure sandboxed execution are essential for effective code RL training. R1-Code-Interpreter [66] introduced a general framework for training LLMs to effectively use a Code Interpreter through multi-stage reinforcement learning. Unlike prior works limited to math or retrieval tasks, it identifies a key challenge that task heterogeneity causes sparse and unstable rewards during RL. To address this, the authors propose a multi-stage curriculum learning approach that prioritizes samples based on their improvement potential. Tool-R1 [67] proposed a sample-efficient reinforcement learning framework that performs multi-step reasoning through executable Python code. It introduces a dynamic sample queue to cache and reuse high-quality trajectories and employs outcome-driven rewards based on code execution success and LLM-judged correctness.

### v3

@@ 4.1 A1: Tool Execution Result as Signal
A1 adaptation treats tool and environment outputs as ground-truth supervision: because the signal is determined by external execution rather than model-internal beliefs, it is verifiable, reproducible, and dense enough to support both supervised and reinforcement-based learning. The key challenge is that such signals optimize individual actions (a single API call, a single code execution) and can miss system-level failures that only surface when multiple actions compose.
@@ 4.1.1 Earlier Works: SFT & Off-Policy Methods
Early A1-type methods train agents from pre-collected data via SFT or DPO. Their shared limitation is an off-policy data distribution: the agent learns from trajectories it did not generate, which can cause a mismatch between training and deployment behavior. Within this family, methods diverge along a key axis—what counts as correct?—progressing from answer-level correctness to format-level validity to raw execution outcomes.
Toolformer [4] (NeurIPS 2023) illustrated the earliest point on this axis: tool outcomes serve as self-supervised signals. The model inserts candidate API calls into text, executes them, and retains a call only if it reduces perplexity beyond a threshold ($L_{i}^{-}-L_{i}^{+}\geq\tau_{f}$). Because the supervision is derived from the model’s own likelihood rather than an external correctness criterion, Toolformer can discover when tools help but cannot distinguish how well a tool call was constructed. Subsequent work addressed this gap through three progressively stronger notions of correctness:
  • Golden-answer alignment: the tool-augmented output must match a known correct response or expert trajectory.
  • Golden-format alignment: the tool call must be structurally and syntactically valid, independent of downstream output.
  • Direct-execution alignment: supervision emerges from actually running the tool, closing the gap between training signal and deployment behavior.
Golden-answer alignment. The simplest instantiation defines correctness as matching a known answer or expert trajectory. A key question within this family is how to exploit negative examples: early methods discard failed trajectories, while later ones treat them as contrastive signal.
ToolAlpaca [53] pioneered the closed-loop generate-execute-evaluate-finetune cycle, where the model calls a tool, observes the runtime outcome (returned values, errors, or completion states), and retrains on the result. Iteration over this loop aligns the model’s internal representation with actual tool semantics and enables generalization to unseen tools—but only positive outcomes contribute to learning. TRICE [54] (NAACL 2024) introduces a ranking loss that scores candidate responses by comparing execution results against ground-truth answers, teaching the model to invoke tools only when they improve accuracy. TP-LLaMA [55] (NeurIPS 2024) goes further by explicitly mining failure information that ToolAlpaca and earlier systems (e.g., ToolLLaMA [10]) discard: at each decision node along a successful trajectory, the expert’s correct next step is preferred ($y_{w}$) and any failed branch is dispreferred ($y_{l}$), and DPO on these pairs converts failure signals into contrastive supervision. The progression from ToolAlpaca to TP-LLaMA thus illustrates a recurring theme in A1 methods: the richer the exploitation of negative feedback, the more data-efficient the adaptation.
Golden-format alignment. A complementary axis defines correctness structurally: a tool call is correct if it is syntactically and logically valid, regardless of whether the final answer is right. This decoupling is valuable when ground-truth answers are expensive to obtain but canonical call formats exist.
Gorilla [56] (NeurIPS 2024) fine-tunes a retrieval-augmented LLaMA model on a large set of machine learning APIs, defining correctness via Abstract Syntax Tree (AST) subtree matching—more robust than string matching because it tolerates differences in parameter order or optional arguments. ToolFlow [57] (NAACL 2025) complements Gorilla by tackling data quality: a tool graph built from parameter and return-value similarities guides the selection of interacting tool subsets, and a planned generation step organizes multi-turn requests for logical consistency. Together, these two works show that format-level supervision can scale to large, evolving API surfaces where golden answers are impractical, but they also reveal a limitation: format correctness does not guarantee functional correctness, so models trained this way can produce well-formed but semantically wrong calls.
Direct-execution alignment. The strongest form of A1 supervision closes the loop entirely: the tool is executed, and its output—success, failure, returned values—becomes the training signal. This eliminates the need for pre-labeled answers or canonical formats but introduces a new challenge: designing execution environments that are safe, reproducible, and informative enough to support learning.
Two design patterns have emerged. The first generates executable code as the action representation. CodeAct [58] (ICML 2024) replaces textual or JSON-based commands with sandboxed code actions, so the environment’s execution feedback directly grounds the model in tool causality. NExT [59] (ICML 2024) applies the same principle to program repair via an iterative Sample-Filter-Train loop: candidate fixes are validated by unit tests, and only passing solutions enter the next fine-tuning round, yielding progressively sharper execution-aware rationales. The second pattern targets autonomous tool discovery. ToolLLM [10] and AutoTools [60] (WWW 2025) have the LLM itself parse raw API documentation into callable functions (tool encapsulation) and then compose multi-function programs to solve user queries (tool programming), with execution outcomes driving iterative self-improvement.
A third pattern applies execution-grounded preference learning to retrieval. LeReT [61] (ICLR 2025) generates diverse search queries, scores retrieved documents via a reward function, and fine-tunes the query generator through Identity Policy Optimization (IPO):
@@
where $\tilde{r}_{\phi}(x,y)=\log\frac{\pi_{\phi}(y|x)}{\pi_{\text{ref}}(y|x)}$ and $\tau$ controls the target margin. Because the reward comes from running the retriever rather than from labeled answers, LeReT adapts to arbitrary off-the-shelf retrievers without modifying the generator. RetPO [62] (NAACL 2025) follows the same logic at lower cost: GPT-4 generates candidate query rewrites, an off-the-shelf retriever (e.g., BM25) scores each by retrieval quality, and a smaller open-source LM is trained via DPO to produce high-reward rewrites.
Summary of SFT & off-policy A1 methods.
The three alignment stages trace a clear arc: from Toolformer’s self-supervised likelihood reduction, through answer- and format-level matching, to direct execution feedback. Each step tightens the coupling between training signal and deployment behavior, reducing the “reality gap” between what the model optimizes and what determines success at inference time. Yet all off-policy methods share a fundamental ceiling: because they learn from pre-collected trajectories, they cannot explore novel tool-use strategies that were absent from the training distribution. This limitation motivates the on-policy RLVR methods described next.
@@ 4.1.2 RLVR-Based Methods
Reinforcement learning with verifiable reward (RLVR) removes the off-policy ceiling of SFT and DPO by letting the agent explore on-line: it proposes actions, executes them, observes verifiable outcomes, and updates its policy—all within the same training loop. The on-policy setting introduces two new design axes absent from the methods above: (i) reward density—whether feedback arrives per step (as in theorem proving) or only at episode end (as in multi-tool reasoning)—which governs credit assignment difficulty; and (ii) reward composition—how task-specific, format, and regularization terms are combined. We organize the domain-specific instantiations below and distill cross-domain principles at the end.
Web search and information retrieval. DeepRetrieval [22] (COLM 2025) established the template: query reformulation is cast as an MDP with retrieval metrics (Recall@K, NDCG, or SQL execution accuracy) as the reward, optimized via KL-regularized PPO:
@@
where $r(q,q^{\prime})=r_{\text{retrieval}}(q,q^{\prime})+r_{\text{format}}(q^{\prime})$. A single framework thereby adapts across literature search, QA retrieval, and text-to-SQL, yielding roughly a threefold recall improvement (65.1% vs. 24.7%) on real-world search engines.
Subsequent work addresses two limitations of this single-step formulation. ReZero [63] adds retry-aware reward shaping via GRPO, teaching the agent to persist after failed searches in partially observable web environments. Orion [64] extends from single-step to multi-turn adaptive search, using turn-level rewards based on normalized similarity and rank; notably, compact 350M–1.2B models learn effective multi-hop search strategies, demonstrating that on-policy RLVR can substitute for large-controller scale when reward density is sufficient.
Code-based tools. Code execution provides a near-ideal A1 environment: feedback is deterministic, sandboxable, and obtainable at every compilation or test-run step. The central design question is how to keep this feedback reliable as tasks grow more heterogeneous. LeDex [65] (NeurIPS 2024) uses a composite PPO reward combining unit-test correctness (CodeBLEU) with explanation quality (semantic similarity), illustrating the value of multi-faceted rewards. RLEF [21] (ICML 2025) formalizes code synthesis as a partially observable MDP where the agent generates, receives public-test feedback, and iterates—showing that multi-turn interaction is key to solving problems beyond the model’s single-shot capability. Code-R1 [66] shifts focus from the policy to the reward pipeline: by eliminating false positives from faulty tests, unsolvable prompts, and mismatched sandboxes, it demonstrates that reward quality dominates reward quantity. R1-Code-Interpreter [67] addresses the heterogeneity of code-interpreter tasks (math, retrieval, data analysis) through multi-stage curriculum learning that prioritizes samples by their improvement potential, mitigating sparse and unstable rewards. Tool-R1 [68] tackles sample efficiency directly via a dynamic sample queue that caches and reuses high-quality trajectories, reducing the exploration cost of multi-step reasoning.